Update for the Week Ending March 4, 2016

Site Remediation Continues

On-going activities:

  • Continued the application of water and calcium chloride for dust control;
  • Continued the application of odor-suppressant foam and BioSolve for dust and odor control;
  • Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site and covered all stockpiles);
  • Continued operation of the WTP;
  • Continued pre-trenching and demolition of subsurface obstructions in the alignment of the Support of Excavation (SOE) wall; and
  • Continued the lining, loading, and shipment of impacted soils to Bayshore Soil Management for treatment and disposal:
    • 18 trucks; 497.17 tons shipped last week; and
    • 819.23 tons shipped to date.

Recently completed activities:

  • Core drilled the Bucket Soil Mix (BSM) In-Situ Stabilization (ISS) demonstration test area and verified that the BSM successfully solidified the material in the Holder No. 2 tank as verified by the Engineer and NYSDEC; and
  • Installed electrical service from Con Edison to the Site and demobilized the temporary generator.

Upcoming activities:

  • Conduct full scale Bucket Soil Mixing (BSM) in the Holder No. 2 tank;
  • Ship removed asphalt for recycling;
  • Pre-drill for deep obstruction removal along the SOE wall alignment;
  • Crush clean, reusable concrete for use as backfill in the deep excavation; and
  • Install dewatering wells and initiate ground water extraction.

Community Air Monitoring Plan Summary

Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.

There were no fifteen-minute Action Limit exceedances of Total Volatile Organics (TVOCs) or odors during this reporting period.

There was one instance of the fifteen-minute average concentrations of Particulate Matter (PM-10) [dust] exceeding the Action Limit on Friday, March 4, 2016, during the delivery of clean slag/cement to the grout plant.  Once the slag/cement was off-loaded from the tanker delivery truck to the Site storage silo, the dust concentration returned to below the Response Limit.  Plastic sheeting will be placed over the delivery truck relief port with future deliveries to mitigate the dust.

If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.

Photos of Activity

Pre-trenching along the SOE wall alignment for the removal of shallow subsurface obstructions.

Loading the lined trucks for shipment of impacted soils to Bayshore Soil Management for treatment and disposal

Cleaning (pressure washing) the trucks on the decontamination pad prior to leaving the Site on North 11th Street

Core drilling the BSM demonstration test area to verify the successful solidification of the material in the Holder No. 2 tank.

Con Edison pulling wire for the electrical service connection to the Site.

Demobilizing the temporary generator from the Site.